OpenStudy (anonymous):

The x-intercept of a line is –5 and the y-intercept of the line is –2. What is the equation of the line?

OpenStudy (anonymous):

A. y = -5/2x -5 B. y = 2/5x + 2 C. y = 5/2x - 2 D. y = -2/5x - 2

OpenStudy (kainui):

Remember that y=mx+b and that b is the y-intercept and m is the slope. What's the slope? Rise over run, so you can consider drawing a picture with the information to figure it out.  now what's the rise over run? start at the left point and you'll see it rises by -2 (since it rises downwards, it's just negative) and runs 5 points to give you a slope of -2/5

OpenStudy (kainui):

Wrong freewilly, you have mixed up your x and y intercepts.

OpenStudy (phoenixfire):

\[y=mx+c\] You know your c=-2. and you know two points. So there's multiple ways to find this out. But an easy way is to just substitute one point into the equation; (-5, 0) \[0=-5m-2=>2=-5m=>{-{2\over 5}}=m\]\[\therefore y=-{2 \over 5}x-2\]

OpenStudy (anonymous):

@Kainui so the answer would be D?

OpenStudy (phoenixfire):

Yes, D is the answer.